13-5 Waymouth Manufacturing operates a contract manufacturing plant located in Dublin, Ireland. The plant provides a...

13-5

Waymouth Manufacturing operates a contract manufacturing plant located in Dublin, Ireland. The plant provides a variety of electronics products and components to consumer goods manufacturers around the world. Cycle time is a critical success factor for Waymouth, which has developed a number of measures of manufacturing speed. Waymouth has studied the matter and found that competitive contract manufacturers have manufacturing cycle times (MCE) of about 32%. When last calculated, Waymouth’s MCE was 25%.

Key measures from the recent month’s production, averaged over all the jobs during that period, are as follows:

Activity Average Time (hours)
New product development 10.0
Materials handling 2.0
Order setup 1.0
Machine maintenance 1.5
Order scheduling 11.0
Inspection of completed order 3.5
Pack and move to storage or ship 1.0
Manufacturing assembly 4.0
Order taking and checking 0.5
Receiving and stocking raw materials 1.0
Inspection of raw materials at start of production 2.5

Required: Determine the manufacturing cycle efficiency (MCE) for the recent month. (Round your answer to 1 decimal place.)
Answer #1

MCE = total processing time/total cycle time= 4/(4+2+1.0+1.5+11.0+3.5+1+1.0+2.5) = 4/27.5 = 14.5%

New product development time and order-taking time are not included in the manufacturing cycle and thus excluded from the cycle time efficiency calculation.